# Optimize the Unseen - Fast NeRF Cleanup with Free Space Prior

### v1

@@ 5 Results
We evaluate our NeRF cleanup method by analyzing performance in the NeRF cleanup task, comparing both quantitative metrics and inference speed. We report numerical results to quantify the reduction of artifacts (floaters) and demonstrate that our approach achieves substantial speedup while maintaining high output quality. Qualitative visualizations further illustrate the effectiveness of our artifact removal process, highlighting that our method not only accelerates the rendering but also preserves or enhances scene fidelity. Additionally, we include ablation studies to assess the impact of key components in our design, offering deeper insights into their contributions to artifact cleanup performance.
@@ 5.1 Dataset and Evaluation Setup
We evaluate the effectiveness of our method in removing artifacts—referred to as ”floaters”—from NeRF reconstructions, following the experimental protocols established by Nerfbusters [26] and BayesRays [8]. Our goal is to assess each method’s ability to eliminate these artifacts while preserving the quality and completeness of the reconstructed scenes. For this evaluation, we utilize the Nerfbusters dataset [26], designed specifically for benchmarking NeRF cleanup methods. This dataset presents challenging scenes for NeRF-based reconstructions, providing an ideal testing ground for artifact removal effectiveness. To gauge performance, we report PSNR and Coverage scores, which quantify reconstruction quality and the extent to which the scene is accurately rendered without artifacts, respectively.
We introduce a key modification to the original experiment to make the task more realistic and challenging. Previous experiments calculated PSNR, SSIM and LPIPS over the ground truth (GT) mask, which only considers the object region. This can overestimate performance, as artifacts outside the GT mask are ignored. In our setup, PSNR is calculated over the predicted mask, penalizing methods that leave artifacts outside the GT mask. This provides a stricter evaluation, ensuring that methods are assessed on their ability to remove artifacts while maintaining scene integrity.
Figure 4: Quantitative Cleanup Results: This plot compares the cleanup performance of various methods using PSNR as a function of Coverage across different thresholds (if applicable). Higher PSNR and Coverage scores represent better performance, with the optimal region located at the upper-right, where both metrics are maximized. For our method, each point reflects the effect of varying cleanup aggressiveness, controlled by the $\lambda$ parameter from Equation 6. Measurements are averaged across all scenes in the Nerfbusters dataset to ensure a comprehensive comparison. We also provide results of Gaussian Splatting for reference.
@@
Implementation Details
All experiments were conducted using the latest version of Nerfacto and Splatfacto from Nerfstudio [25], pre-trained for 30,000 steps. Comparisons between different methods were performed on the same pre-trained base NeRF model to ensure consistency. In our method, density optimization was achieved with 1,000 iterations using $2^{17}$ randomly sampled points across the 3D scene space. Optimization was carried out through Nerfacto’s optimizers. All experiments were run on a single NVIDIA RTX A5000 GPU. Results for BayesRays [8] and Nerfbusters [26] were generated using their respective pipelines, including the pre-trained 3D diffusion model weights provided by Nerfbusters. The Nerfacto model we used incorporates recent advancements in NeRFs, including hash grid encoding [18], proposal sampling and scene contraction [1], as well as per-image appearance optimization [14]. The splatfacto model is based on Gaussian Splatting [11] with regularization to prevent artifacts [29, 27].
Figure 5: Runtime Comparison: Training runtime Vs Inference runtime per frame. This plot compares the runtime performance of three cleanup methods. The inference time is measured in milliseconds per frame, averaged across all scenes in the Nerfbusters dataset. Note that both Nerfbusters and Free Space Prior methods only involve the finetuned NeRF during inference, resulting in the same inference time for both.
@@ 5.2 Performance Comparison
In this section, we assess the effectiveness of our NeRF cleanup method in comparison to state-of-the-art approaches, focusing on both artifact removal quality and computational efficiency. We evaluate performance through quantitative metrics such as PSNR and Coverage, as well as runtime efficiency, considering both training and inference times. Our method is designed to maximize the trade-off between image quality and coverage while providing significant speed advantages.
@@
Our method achieves competitive results with significantly faster training and inference times, as shown in Figure 5. In Figure 4, our approach demonstrates high coverage while maintaining a PSNR comparable to BayesRays, the current state-of-the-art in NeRF cleanup. Striking a balance between high PSNR and high coverage is crucial, as high PSNR with low coverage can be misleading, as illustrated, as shown in Figure 3. For instance, BayesRays at a 0.3 threshold achieves the highest PSNR on the graph but suffers from low coverage, leading to noticeable gaps in the reconstructed scene that compromise overall image quality. In contrast, both the Nerfacto baseline and Gaussian Splatting models exhibit excellent coverage but lower PSNR, underscoring the challenge posed by the dataset, which is designed to test these specific issues. Artifacts such as density accumulation in Nerfacto and scattered Gaussians in Gaussian Splatting are also clearly visible.
Our method, as well as BayesRays with a 0.9 threshold, produces visually superior results that retain the scene’s structure and coverage, making them more visually appealing and accurate to a human eye. These images highlight that our method preserves both high coverage and high PSNR, outperforming Nerfbusters, which, while comparable in inference, achieves lower PSNR and coverage than our method. Overall, our method achieves the best coverage, surpassed only by the baseline, but with substantially higher PSNR, confirming its effectiveness in both artifact removal and scene fidelity.
Figure 6: Density Histogram: Histogram of densities across the Pikachu scene after sigmoid-softened as in Equation 4, shifted to [0, 1] (Top) Histogram of densities across the scene before the cleanup process (Bottom) Histogram of the densities across the scene after the cleanup process.
@@
Our method combines fast training and inference times Figure 5 while maintaining high-quality artifact cleanup Figure 3. For inference, both Nerfbusters and our method achieve a per-frame rendering time of 205 ms, as each directly fine-tunes the NeRF without adding additional overhead at inference. In contrast, BayesRays incurs a significant slowdown, with a per-frame inference time of 506 ms—2.5 times slower—due to its additional uncertainty thresholding applied during rendering.
@@
For training time, Nerfbusters employs a pre-trained 3D diffusion model to fine-tune NeRF per scene, resulting in a high training time of approximately 20 minutes per scene. BayesRays is faster, taking 37 seconds to gather data for Hessian computation, while our approach is the fastest, completing cleanup training in just 25 seconds.
Figure 7: Density Along a Ray: (Left) Novel views of the Plant scene, rendered with the Nerfacto Baseline and our Free Space Prior, with a red dot marking the sampled ray shown in the graph. (Right) Graph of density along the marked ray. Before cleanup, the density field is noisy, while after cleanup it captures only the surface along the ray. In the baseline densities, the first peak corresponds to the floater seen in the image, the second peak to the wall (present in both images and graphs), and the subsequent unseen regions, which are noisy in the baseline, are enforced to zero after cleanup. Densities are softened using Equation 4 and shifted to [0, 1]
Overall Performance
Our NeRF cleanup method demonstrates a well-rounded balance of high-quality artifact removal, high scene coverage, and rapid computational performance. Compared to existing methods, it achieves similar or superior visual results with notably reduced training times and maintains efficient inference speeds on par with the baseline NeRF model. These results underscore the practicality of our approach for applications where both visual fidelity and speed are critical. The combination of efficient cleanup and low runtime costs illustrates our method’s capability to enhance NeRF-based scene reconstructions without the need for extensive additional resources.
@@ 5.3 Ablation Study
First, we examine the effect of our method on the density distribution in the scene, comparing density histograms before and after cleanup. This analysis helps us understand how our approach distinguishes between object surfaces and empty space, ultimately refining the NeRF’s representation of the scene. Second, we test our hypothesis that sampling across the entire 3D space, especially in unseen regions, is crucial for effective artifact removal. We conduct experiments to compare our random sampling strategy with sampling exclusively along training rays, evaluating the effect on cleanup quality. Together, these studies provide insights into why our method succeeds in both comprehensive artifact removal and scene preservation.
Densities Histogram
We analyzed density distributions across the scene before and after applying our cleanup method, by measuring the densities in random $2^{18}$ points across the scene space, visualized in Figure 6. Before cleanup, the density histogram shows three distinct ranges: high densities representing object surfaces and floaters, intermediate densities indicating regions with inconsistent optimization from the NeRF training, and zero densities corresponding to empty space. The presence of intermediate densities highlights that regions outside the training camera views received little to no regularization during the NeRF training, allowing densities to vary without specific constraints.
After cleanup, the density distribution shifts to a bimodal pattern, effectively removing intermediate values and clustering densities around zero or high values. This shift indicates that our method enforces a clear distinction between object and empty space, aligning the density distribution more closely with an ideal NeRF scene representation. The majority of the densities now approach zero, with only a small fraction remaining high to capture object surfaces. This refined distribution improves the NeRF’s alignment with real scene structures, as illustrated in Figure 1. As shown in Figure 7, the post-cleanup distribution clearly demonstrates this shift, with the intermediate densities largely eliminated and a more concentrated focus on the object and empty space densities.
Sampling Along Training Rays
In our approach, we apply the Free Space Prior Loss Equation 5 by sampling randomly across the entire 3D scene space, which includes both seen and unseen regions. We hypothesize that sampling in unseen areas is crucial to effectively removing artifacts and achieving high-quality novel views. To validate this, we conducted an ablation study comparing our method with a setup in which the Loss is applied only to densities sampled along the training rays. These rays pass through seen regions and some unseen areas directly behind objects, but do not cover all unseen areas (as illustrated in Figure 2).
@@
Figure 8: Applying the Prior Along Training Rays: Ablation study on applying the Free Space Prior loss specifically along the training rays, with varying numbers of samples per ray. (Top) Visualizations show the cleanup results for each method on a novel view. (Bottom) The PSNR score for each method, averaged over the Pikachu scene evaluation set, demonstrate the quantitative impact of sampling along training rays.
In this study, we tested two variations limited to ray-based sampling. The first used 32 samples per ray across 4,096 rays, matching the 131,072 samples of our original method. The second increased samples to 512 per ray, totaling 2,097,152 points per iteration. Figure 8 show that our random sampling achieves a PSNR of 25.6, significantly surpassing ray-based methods. Sampling with 32 samples per ray raised PSNR from 19.3 to 19.9, while 512 samples per ray improved it to 20.6, though artifacts persisted, and some scene details were lost due to over-aggressive cleanup.
These findings confirm that sampling the full 3D space is essential for effective artifact removal, as ray-based sampling misses unseen regions, resulting in incomplete cleanup or loss of scene detail.
@@ 6 Conclusion
we present an efficient post-hoc cleanup method for Neural Radiance Fields (NeRFs) that is based on a simple Free Space Prior, and show that it significantly reduces artifacts, specifically floaters, in unseen regions of a scene. Our approach stands out due to its ability to eliminate these artifacts without modifying the core NeRF architecture or adding additional memory overhead, offering a solution that is both fast and scalable. We have demonstrated that our method outperforms state-of-the-art approaches in terms of both efficiency and quality. Specifically, our method achieves up to 2.5×2.5$\times$2.5 × faster inference time, 1.5×1.5$\times$1.5 × faster train time and requires no additional memory. It maintains high fidelity in the reconstructed scene, making it a practical choice for real-world applications.
7 Supplementary
7.1 Method Robustness

### v3

@@ 5 Results
Table 1: Runtime Comparison: Inference time (ms/frame) is averaged across all Nerfbusters scenes. Both Nerfbusters and Free Space Prior use only the fine-tuned NeRF at inference, resulting in identical inference runtime.
Method Cleanup Time (s) Inference Time (ms)
Nerfbusters 1200 205
BayesRays 37 506
Free Space Prior (Ours) 25 205
We evaluate our NeRF cleanup method by analyzing its performance in the NeRF cleanup task, comparing both quantitative metrics and computational efficiency in terms of cleanup and inference speed. We report numerical results to quantify the reduction of artifacts (floaters) and demonstrate that our approach achieves significant speedup while maintaining high-quality novel view synthesis.
Beyond numerical evaluations, we qualitatively assess our method across a diverse range of NeRF architectures and datasets, showcasing its effectiveness in different scenarios. Additionally, we conduct ablation studies to analyze the impact of key components in our design, offering insights into their role in artifact removal. The implementation details are in the supplementary Sec.˜A.2.
@@ 5.1 Dataset and Evaluation Setup
We follow the experimental protocols established by Nerfbusters [31] and BayesRays [10]. Our goal is to assess each method’s ability to eliminate these artifacts while preserving the quality and completeness of the reconstructed scenes. For quantitative evaluation, we use the Nerfbusters dataset [31], which is the standard benchmark for NeRF cleanup methods. Unlike other datasets, which primarily focus on reconstructing regions well-covered by training views, Nerfbusters uniquely provides evaluation views that are significantly farther from the training cameras, making it the most challenging and relevant dataset for artifact removal. This ensures that cleanup methods are tested under realistic conditions where unseen regions contribute to artifacts. In addition to the quantitative evaluation on Nerfbusters, we qualitatively evaluate our method across a diverse range of NeRF architectures and datasets, demonstrating its generalizability and effectiveness in different scenarios.
Figure 5: Quantitative Cleanup Results: PSNR vs. Coverage comparison of cleanup methods across different thresholds / $\lambda$. Higher values indicate better performance, with the optimal region in the upper-right corner. Results are averaged across all Nerfbusters scenes for a comprehensive evaluation.
@@
Baseline
Our baseline model (Nerfacto) is built on modern NeRF architectures, incorporating several recent advancements, including hash grid encoding from iNGP [23], proposal sampling and scene contraction from MipNeRF[1], distortion loss from MipNeRF360 [2] as well as per-image appearance optimization from NeRF-in-the-Wild [19]. While these techniques provide a degree of regularization, they fail to prevent the accumulation of artifacts in unseen regions. Our Free Space Prior directly addresses this issue by enforcing explicit density constraints in these areas.
To measure performance, we report PSNR and Coverage scores, which quantify both reconstruction quality and the extent to which the scene is accurately rendered without artifacts.
In addition to this standard evaluation (Sec.˜A.4), we introduce and evaluate a modified experiment that makes the task more realistic and challenging: previous evaluations calculated PSNR, SSIM, and LPIPS using the ground truth (GT) mask, which only considers the object region. This can overestimate performance, as artifacts outside the GT mask are ignored. In our setup, the metrics computed over the predicted mask, penalizing methods that leave artifacts outside the GT mask (Sec.˜A.6). This stricter evaluation ensures that methods are assessed on their true ability to remove artifacts while preserving scene integrity.
@@ 5.2 Performance Comparison
We evaluate our approach both qualitatively (Fig.˜3) and quantitatively (Fig.˜5), using the challenging modified evaluation (i.e., using the predicted mask to evaluate results).
Striking a balance between PSNR and coverage is crucial, as high PSNR with low coverage can be misleading. For instance, as shown in Fig.˜5, BayesRays at a 0.3 threshold achieves the highest PSNR but suffers from low coverage, leading to noticeable gaps in the reconstructed scene (Fig.˜3) that compromise overall image quality.
As also seen in Fig.˜5, Free Space Prior demonstrates high coverage while maintaining a PSNR comparable to BayesRays, the current state-of-the-art in NeRF cleanup. Among the other methods, Nerfbusters exhibits significantly lower coverage and PSNR, indicating excessive artifact removal. In contrast, L2 regularization, Gaussian Splatting, and the Nerfacto baseline exhibit excellent coverage but lower PSNR, retaining more scene information while struggling with artifact suppression and overall scene fidelity, highlighting the dataset’s challenge.
Our method, as well as BayesRays with a 0.9 threshold, produce visually superior results that retain the scene’s structure and coverage, making them more visually appealing and accurate to a human eye. These images highlight that our method preserves both high coverage and high PSNR, outperforming Nerfbusters and L2 regularization, which, while comparable in inference time, achieves lower PSNR and coverage than our method.
For completeness, we show that our method also works if applied during the training phase when learning the NeRF from scratch. In that case, our method makes the cleanup phase redundant. The visual results ($\lambda=10^{-5}$) are shown in Fig.˜3 and the numerical details are in the appendix Tab.˜4.
@@
Our method combines fast cleanup and inference times (Tab.˜1) while maintaining high-quality artifact cleanup (Fig.˜5). For inference, both Nerfbusters and our method achieve a per-frame rendering time of 205 ms, as each directly fine-tunes the NeRF without adding additional overhead at inference. In contrast, BayesRays incurs a significant slowdown, with a per-frame inference time of 506 ms—2.5 times slower—due to its additional uncertainty thresholding applied during rendering.
@@
For cleanup time, Nerfbusters employs a pre-trained 3D diffusion model to fine-tune NeRF per scene, resulting in a high cleanup time of approximately 20 minutes per scene. BayesRays is faster, taking 37 seconds to gather data for Hessian computation, while our approach is the fastest, completing cleanup training in just 25 seconds.
Figure 6: Density Histogram: Histogram of densities across the Pikachu scene after sigmoid-softened, shifted to [0, 1]. (Left) Histogram of densities across the scene before the cleanup process. (Right) Histogram of the densities across the scene after the cleanup process.
Table 2: Comparison of methods on geometric metrics.
Method Chamfer Distance Depth (RMSE) Depth (MAE) PSNR
Ours () 0.011 3.80 1.02 28.64
BayesRays (T=0.9) 0.011 3.91 1.07 28.51
5.3 Geometric Evaluation
While NeRF is a photometric based method, it represents a 3D scene with geometric properties. To evaluate the cleanup quality, in addition to the photometric evaluation we conducted three geometric experiments on the Nerfbusters dataset as shown in Table˜2.
Chamfer Distance: We measured Chamfer Distance between pre-cleanup and post-cleanup scenes. For each scene, we generated 35k-point point clouds using training cameras and NeRF reconstructions (pre- and post-cleanup), then computed Chamfer Distance between them.
Depth Consistency: We evaluated surface preservation using depth maps to check whether they remain consistent after cleanup on the training images. For the most accurate measurement, we compared rendered depth maps to pseudo-gt depth maps derived from a NeRF trained on both training and test images.
PSNR on Training Images: We measured the PSNR of training rendered images after cleanup. The rationale is that validating scene integrity using training images ensures that the core mechanism of NeRF reconstruction is preserved even through the cleanup phase.
All three geometric evaluations confirm that our uniform sampling strategy preserves geometric details effectively as our method achieves comparable or better scores compared to BayesRays.
5.4 Robustness
Our method is designed to be robust across different NeRF architectures and datasets, ensuring effective artifact removal across a variety of conditions. Unlike methods that rely on architecture-specific modifications or learned priors trained on specific datasets, our Free Space Prior provides a generalizable solution that applies seamlessly to different NeRF formulations.
As shown in Fig.˜1, we evaluate our method across diverse NeRF architectures, including Nerfacto [30], Instant-NGP [23], LeRF [14], Seathru-NeRF [16], and Open-NeRF [8], each of which has distinct optimization techniques and applications. Our approach also generalizes well across challenging datasets, such as Nerfbusters [31], MipNeRF-360 [2], LeRF [14], Seathru [16], and KITTI-360 [17], covering a range of real-world and synthetic environments. These results highlight the versatility of our method, effectively reducing artifacts without requiring architecture-specific tuning.
Our prior works on the Seathru-NeRF dataset, which operates in extreme underwater conditions. The physics of underwater imaging involves attenuation and back scatter, so density should not be zero. Yet, our prior reduces artifacts significantly. See Fig.˜1, column 4. Beyond novel view synthesis, our cleanup process enhances downstream tasks such as NeRF-based semantic segmentation and language-based scene understanding, as shown in Fig.˜1, columns 3 and 5.
5.5 Method Analysis
To better understand the effect of our Free Space Prior, we analyze how it modifies the density distribution within the NeRF.
Density Histogram Analysis
We visualize the density distribution before and after cleanup by sampling points randomly across the 3D space and plotting their density values, as shown in Fig.˜6. Before cleanup, the density histogram exhibits a wide distribution, with intermediate densities appearing throughout the scene. These intermediate values correspond to unwanted density accumulations (floaters) in unseen regions, which lack sufficient supervision during NeRF optimization.
After applying our Free Space Prior, the density distribution shifts toward a bimodal structure, where densities are either near zero (free space) or high (scene surfaces). This indicates that our method successfully removes floaters while preserving scene structures, leading to a cleaner and more physically meaningful density field. See Sec.˜A.9 for underwater analysis.
Density Along a Ray
To further illustrate the effect of our cleanup, we examine the density values along a sampled ray, as shown in Fig.˜4. Before cleanup, the ray passes through multiple unintended density peaks, corresponding to floating artifacts that degrade novel view synthesis. After cleanup, these spurious densities are eliminated, and the ray exhibits only a single, well-defined peak at the true surface, confirming that our method effectively enforces physically consistent density distributions.
Figure 7: Number of Samples Study: PSNR vs. Cleanup Time of the Pikachu scene for different NN values. Our choice of $N=2^{17}$ marked in red.
@@ 5.6 Ablation Study
The Number of Randomly Sampled Points
The rational behind selecting $N$ is that $N$ should be proportional to the reconstruction loss sampling density. In each iteration, we use 4,096 rays with 352 samples per ray ( 1.45M samples) for reconstruction loss. We chose ( 131K samples), which is approximately one order of magnitude smaller, ensuring the FSP loss provides a meaningful cleanup signal while adding minimal computational overhead.
The ablation study in Fig.˜7 confirms our rationale, $N=2^{17}$ achieves substantial cleanup improvements with only 10% additional runtime compared to minimal sampling, while avoiding the diminishing returns of higher values.
Table 3: Comparison of different methods under sparse and dense view settings.
Sparse View Dense View
# Train Images 24 images (Every 8th) 47 images (Every 4th) 93 images (Every 2nd) 161 images (Regular) 8 Out of 9
Method PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
Base 22.32 0.637 0.261 24.41 0.703 0.226 25.33 0.724 0.212 25.74 0.731 0.212 27.74 0.789 0.181
BayesRays 22.33 0.637 0.259 24.42 0.703 0.226 25.38 0.725 0.212 25.74 0.731 0.212 27.73 0.789 0.181
Ours 22.31 0.640 0.257 24.40 0.705 0.222 25.35 0.727 0.211 25.74 0.732 0.213 27.70 0.788 0.183
Sparse and Dense View Performance
To overcome concerns around performance degradation we conducted an ablation study with sparse and dense view settings, comparing our method with $\lambda=0.01$ and BayesRays with $T=0.9$, which achieve comparable coverage rates. The sparse view study is conducted on the Garden scene from mipNeRF-360 dataset. The results on Table˜3 demonstrate that our method not only preserves NeRF’s interpolation capabilities under sparse training conditions but actually improves perceptual quality (SSIM/LPIPS) while maintaining comparable PSNR.
The dense view study is conducted on the mipNeRF-360 dataset and as shown in Table˜3 our method has a negligible influence and it maintains high-quality results while being 2.5× faster in inference compared to BayesRays.
@@
Figure 8: Enforcing the Prior Along Training Rays: (Top) cleanup result. (Bottom) PSNR on Pikachu scene.
Forcing Prior Along a Ray
A key design choice in our method is sampling density across the entire 3D space. We conduct an ablation study to assess the impact of this global sampling strategy on artifact removal.
An alternative approach would be to apply the prior only to points sampled along training rays, as commonly done in NeRF-based regularization methods. We compare our full 3D sampling strategy with one that applies the prior loss solely along training rays, varying the number of samples per ray. As shown in Fig.˜8, limiting optimization to training rays results in incomplete cleanup and leaves floaters in unseen regions. Even with more samples per ray, cleanup improves only slightly and does not match the performance of full 3D sampling, which better regularizes completely unseen regions.
These results confirm that sampling beyond training rays is essential for effective NeRF cleanup. By enforcing the Free Space Prior across the entire 3D scene, our method ensures that unseen regions remain artifact-free, leading to significantly improved novel view quality.
6 Limitations
The method optimizes unseen regions while using reconstruction loss to preserve the original scene. As a result, it cannot correct artifacts caused by image inconsistencies, moving objects, poor bundle adjustment, or noisy input images. It is only effective for artifacts resulting from under-optimization in specific regions of the scene. An example of the limitation using noisy images is in Sec.˜A.8.
@@ 7 Conclusion
We present an efficient cleanup method for NeRFs that leverages a simple yet effective Free Space Prior to significantly reduce artifacts, particularly floaters, in unseen regions. Our approach eliminates these artifacts without modifying the NeRF architecture or increasing memory overhead, making it a lightweight and scalable solution. We demonstrate that our method outperforms SOTA cleanup approaches in both efficiency and quality, achieving up to $\mathbf{2.5\times}$ faster inference, $\mathbf{1.5\times}$ faster cleanup, and requiring no additional memory. Furthermore, our approach is robust across diverse NeRF architectures and datasets, effectively reducing artifacts in a variety of real-world and synthetic scenarios. By preserving scene fidelity while ensuring consistent and generalizable artifact removal, our method provides a practical and widely applicable solution for enhancing NeRF-based NVS.
Appendix A Technical Appendices and Supplementary Material
A.1 Appendix Overview
This document contains supplemental material regarding the following topics:
  • A.1 Implementation Details
  • A.2 Appendix Overview
  • A.3 Method Robustness - Discussion and visualization of the robustness of our method.
  • A.4 Evaluation Using the Old Paradigm - Evaluation using the original Nerfbusters evaluation.
  • A.5 Choosing $\lambda$ - Discussion about the temperature parameter $\lambda$
  • A.6 Full Quantitative Results - The numerical results including a discussion about Dice score as an alternative to coverage metric.
  • A.7 Additional Qualitative Results - more visual examples of our method.
  • A.8 Noisy images Study - Study about the limitations.
  • A.9 Densities in Complex Medium - Underwater analysis.
A.2 Implementation Details
All experiments were conducted using the latest version of Nerfacto and Splatfacto from Nerfstudio [30], pre-trained for 30,000 steps. Comparisons between different methods were performed on the same pre-trained base NeRF model to ensure consistency. In our method, density optimization was achieved with 1,000 iterations using $2^{17}$ randomly sampled points across the 3D scene space. Optimization was carried out through Nerfacto’s optimizers. All experiments were run on a single NVIDIA RTX A5000 GPU. Results for BayesRays [10] and Nerfbusters [31] were generated using their respective pipelines, including the pre-trained 3D diffusion model weights provided by Nerfbusters. The predicted mask indicates the presence of "something" in a pixel after cleanup, created by thresholding the density accumulation at 0.98. The splatfacto model is based on Gaussian Splatting [13] with regularization to prevent artifacts [34, 32].
A.3 Method Robustness